OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 3x + y = 10 y = x - 2 A. (4, 2) B. (3, 1) C. (2, 0) D. (5, 3)

OpenStudy (diyadiya):

substitute y in equation 3x + y = 10 3x+(x-2)=10

OpenStudy (diyadiya):

3x+x-2=10 4x-2=10 add 2 on both sides and divide by 4 what's x now ?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 3x + (x - 2) = 10 }\) \(\Large \color{purple}{\rightarrow 3x + x - 2 = 10 }\) \(\Large \color{purple}{\rightarrow 3x + x = 12 }\) \(\Large \color{purple}{\rightarrow 4x = 12 }\) \(\Large \color{purple}{\rightarrow x = 3 }\) y = x - 2. Clearly, \(\Large \color{purple}{\rightarrow y = 3 - 2 = 1 }\)

Parth (parthkohli):

@daijagattison btw, did you understand my tutorial?

OpenStudy (anonymous):

YESS i understand it very well you doing a god job helping out